# The role of quantum and classical correlations in shrinking algorithms for optimization

### v1

@@ Abstract
The benefit of quantum computing for solving combinatorial optimization problems (COPs) constitutes an open research question. In this work, we study the performance of a shrinking algorithm for COPs. The algorithm leverages correlations extracted from quantum or classical subroutines to recursively simplify the problem. We compare the performance of the algorithm equipped with correlations from the quantum approximate optimization algorithm (QAOA) as well as the classical linear programming (LP) and semi-definite programming (SDP) relaxations. This allows us to benchmark the utility of QAOA correlations against established classical relaxation algorithms. We apply the recursive algorithm to MaxCut  problem instances with up to a hundred vertices at different graph densities. Our results indicate that LP outperforms all other approaches for low-density instances, while SDP excels for high-density problems. Moreover, the shrinking algorithm proves to be a viable alternative to established methods of rounding LP and SDP relaxations. In addition, the recursive shrinking algorithm outperforms its bare counterparts for all three types of correlations, i.e., LP with spanning tree rounding, the Goemans-Williamson algorithm, and conventional QAOA. While the lowest depth QAOA consistently yields worse results than the SDP, our tensor network experiments show that the performance increases significantly for deeper QAOA circuits.
@@ I Introduction
Combinatorial optimization problems (COPs) are highly relevant in diverse sectors across both science and industry. However, they are in general challenging to solve efficiently because of the exponential growth of the solution space with increasing problem size. Traditionally, a variety of classical strategies, including linear programming (LP) [1] and semi-definite programming (SDP) relaxations [2], have been applied to find approximate solutions to COPs. Recently, approaches utilizing quantum resources have emerged as promising alternative solution strategies [3]. Among these, the quantum approximate optimization algorithm (QAOA) [4] is of particular interest because of its universality and adaptive complexity. QAOA is a hybrid quantum-classical variational algorithm designed to find approximate solutions to unconstrained binary optimization problems using noisy-intermediate scale quantum (NISQ) devices and beyond.
@@
It is unclear whether NISQ algorithms can in practice provide speedups over classical methods for solving COPs. In fact, several issues of the near-term variational approaches have been identified, such as barren plateaus [5] and noise of quantum devices [6]. Another issue of QAOA is its locality, i.e., the effect that only qubits that are separated by less than a certain distance in the graph representation of a problem can interact with each other at a given circuit depth. This property can be used to prove limitation results on the performance of QAOA [7, 8, 9, 10].
@@
The recursive QAOA (RQAOA) was introduced by Bravyi et al. [7, 11] to overcome the locality-induced limitations of QAOA. The algorithm operates by recursively simplifying the problem based on correlations between variables obtained from QAOA. The two steps of computing the correlations and fixing the variables are executed iteratively until the problem is fully solved. By introducing new connections between previously unlinked variables, RQAOA overcomes the locality of QAOA. Several works have generalized and extended RQAOA. These proposals include using problem-specific update rules [12, 13], analog quantum devices [12] or a shrinking procedure based on the classical calculation of correlations [14]. It has been shown that those recursive algorithms outperform the original QAOA for many problem instances [15, 11, 16]. However, it is not clear whether the good results are due to the recursive shrinking procedure or the quantum correlations. Here, we aim to give insights in this direction.
@@
This study focuses on evaluating how different classical and quantum methods for calculating correlations affect the performance of a shrinking algorithm similar to RQAOA. The investigated routines for calculating correlations are the classical LP and SDP relaxations and QAOA. Our numerical experiments on instances of MaxCut allow us to benchmark the utility of quantum correlations against correlations obtained from established classical approximation algorithms. Moreover, we find that the shrinking procedure employed here is a viable alternative to traditional rounding routines for LP and SDP relaxations.
@@
The remainder of the paper is organized as follows. In Section II we introduce the MaxCut problem and the different routines (LP, SDP and QAOA) for calculating the correlations. In addition, we present the shrinking algorithm and how it utilizes the calculated correlations. In Section III we analyze the performance of the shrinking algorithm employing different correlation sources, before discussing the implications of our findings in Section IV. In Section V we conclude by suggesting potential future research directions.
@@ II Methods
We begin this section by introducing two mathematical formulations of the MaxCut  problem. Next, we provide details on how the correlations can be computed using the LP, SDP, and QAOA routines. Finally, we present the shrinking algorithm that uses the correlations to recursively simplify the optimization problem.
@@ II-A MaxCut problem and its encoding
Despite the simple structure of the MaxCut problem, it constitutes a popular example of a COP since any quadratic unconstrained binary optimization (QUBO) problem can be transformed into a MaxCut problem [17, 18, 19, 20]. Furthermore, solving it is NP-hard [21], and finding solutions for dense instances with hundreds of variables can already overstrain state-of-the-art algorithms [22].
A MaxCut problem instance is defined by a weighted undirected graph $G=(V,E)$ with vertices $V=\{i\}$, edges $E=\{e\}$ and edge weights $w_{e}$. We denote the number of vertices by $n=|V|$. In MaxCut, we are tasked with finding a subset of nodes that maximizes the weight of edges connecting the chosen node subset and its complement. Formally, we want to find a node partition $W\subseteq V$ such that the edge set $\delta(W)\coloneqq\{ij\in E\mid i\in W,j\in V\setminus W\}$ maximizes its weight defined as $\sum_{e\in\delta(W)}w_{e}$.
The weighted MaxCut problem can also be formulated as maximizing an integer quadratic unconstrained cost function $C(\mathbf{x})$ in the form of
@@ II-B Means of computing correlations
At the core of the algorithm are the correlations between decision variables of a MaxCut problem. Each edge $ij\in E$ in the problem graph is assigned a correlation $b_{ij}\in[-1,1]$. Ideally, we want this correlation to be indicative of the correlation between variables in high-quality solutions of the MaxCut problem. In this case, a large negative (positive) correlation between two variables indicates that in good candidate solutions, these two variables take mostly opposite (equal) values on average. In the language of MaxCut, this translates to an edge predominantly being cut for negative correlations and not cut for positive correlations.
The shrinking algorithm works identically irrespective of the correlations that are used. However, for different situations, different means of computing correlations might be better suited to achieve the best possible results. Thus, in this paper, we compare three means of computing correlations, namely, using LP and SDP relaxations, and QAOA.
@@ II-B1 Linear programming (LP)
For a weighted undirected graph $G=(V,E)$, we define an edge-incidence vector $\mathbf{y}\in\{0,1\}^{|E|}$. Here, 00 corresponds to the edge not being cut, and $1$ represents the edge being cut.
An integer linear programming formulation of the MaxCut problem can be weitten as (cf. [23, 24, 25, 26])
@@
$\begin{split}&y_{e}\in\{0,1\},\ \forall e\in E,\end{split}$ (2d)
where equations (2b) are the so-called odd-cycle inequalities, known to be sufficient to define a cut. While the MaxCut  problem defined by Eqs. (2a), (2b), (LABEL:eq:_odd-cycle_d) cannot be solved efficiently in general, the linear relaxation defined by Eqs. (2a)–(LABEL:eq:_odd-cycle_c) can be solved in polynomial time, e.g., by a branch-and-bound approach, and yields an upper bound on the maximum cut [22]. This also means that if the relaxation returns an integer solution, the problem has immediately been solved to optimality by the relaxation.
Analogously to Ref. [14], we compute correlations from an optimum solution $\tilde{\mathbf{y}}\in[0,1]^{|E|}$ to the linear program Eqs. (2a)–(LABEL:eq:_odd-cycle_c) by the means of the affine function
$b_{e}^{\mathrm{LP}}\coloneqq 1-2\tilde{y}_{e}\in[-1,1].$ (3)
The intuition behind these correlations is that we get strong correlations when either the edge weight is large or when the decision to cut an edge does not interfere with other desired cuts. Here, the term interfering stands for the following situation: Say, we are deciding to cut a specific edge. This decision makes it impossible to cut other edges that we might want to cut as well. In such a case we assign a worse correlation to this specific edge to circumvent cutting it.
@@ II-B2 Semi-definite programming (SDP)
The calculation of semi-definite programming correlations is inspired by the seminal approximation algorithm introduced by Goemans and Williamson [2], which is the best known polynomial-time approximation algorithm for MaxCut  problems. The main idea underlying the Goemans-Williamson (GW) algorithm is that we replace the NP-hard integer quadratic problem formulation of MaxCut as introduced in Eq. (1) by a relaxed version of the problem. The relaxation admits a larger solution space that contains all possible solutions to the integer problem. Thus, the optimal solution of the relaxed problem upper bounds the best solution of the integer model by design. Concretely, we replace integer values with multi-dimensional vectors $\mathbf{v}_{i}\in\mathbb{R}^{n}$ with $\norm{\mathbf{v}_{i}}=1$, i.e., to the $(n-1)$-dimensional unit sphere $S_{n-1}$. Using this, we can relax the maximization of Eq. (1) to
@@
where $L\coloneqq\text{diag}(A\mathbf{{e}})-A$ is the Laplacian of the problem graph and $\mathbf{{e}}$ stands for the vector of all ones. Here, $\langle a,b\rangle=\text{tr}(ab^{\text{T}})$ represents the Frobenius inner product of two matrices $a$, $b\in\mathbb{R}^{n\times m}$. This generic semi-definite program can be solved in polynomial time by various solvers such as cvxopt, see, e.g., [27, 28, 29].
@@
From the vectors $\{\mathbf{v}_{i}\}$, the correlation for each edge $\{i,j\}\in E$ is obtained by evaluating the dot product
@@
which measures the level of (anti-)alignment of the vectors. This way we assign a large absolute correlation if the vectors are approximately parallelly or anti-parallelly aligned, i.e. the dot product takes a value close to $+1$ or $-1$. Comparing Eqs. (1) and (4) reveals that this relaxed formulation is analogous to the case of integer variables, where an edge is cut if $x_{i}x_{j}=-1$, and not cut for $x_{i}x_{j}=+1$. In contrast to the integer formulation, the relaxed dot product correlations can now take values between $-1$ and $+1$. Furthermore, we can compute these correlations efficiently.
@@
Alternatively, we can obtain correlations by leveraging the hyperplane rounding procedure from the GW algorithm [2]. The rounding procedure maps vectors to an integer solution, yielding an approximation of the maximum cut. This is done by first choosing a random hyperplane, defined by a normal vector $\mathbf{r}\in S_{n-1}$. Next, we assign the vertices $V$ into two partitions $S\subseteq V$ and $\overline{S}=V\,\backslash\,S$ using the criterion:
$\quad S=\{i\ |\ \mathbf{v}_{i}\cdot\mathbf{r}\geq 0\}.$ (7)
As the evaluation of a given candidate cut can be performed efficiently, it is possible to try several random hyperplanes, calculate the resulting cut value, and choose the hyperplane corresponding to the best cut value. It has been shown in Ref. [2] that the cuts obtained this way are guaranteed to achieve an approximation ratio of at least $0.878$.
@@
The GW correlations are designed to combine both the global information about the assignment of the vectors to one of the two partitions defined by the hyperplane, as well as the local information about the (anti-)parallel alignment of the two vectors.
@@ II-B3 QAOA
Finally, we turn to computing the correlations using QAOA. The goal of QAOA is to find an $\mathbf{x}\in\{-1,1\}^{n}$ that maximizes the cost function $C:\{-1,1\}^{n}\rightarrow\mathbb{R}$ of an arbitrary integer quadratic unconstrained optimization problem. Starting from the uniform superposition state $\ket{+}$ defined by
@@
is maximized. Thus, we hope to prepare a superposition of high quality candidate solutions.
As follows from Eq. (1), for MaxCut the cost Hamiltonian $H_{C}$ can be written as
@@
where $I$ is the identity operator and $Z_{i}$ is the Pauli-Z operator acting on qubit $i$. After maximizing the expectation value $F(\boldsymbol{\beta},\boldsymbol{\gamma})$, correlations $b_{ij}$ between nodes $i$ and $j$ connected by an edge can be defined as the expectation value with respect to the QAOA state:
@@
where $\boldsymbol{\beta_{\text{opt}}}$ and $\boldsymbol{\gamma_{\text{opt}}}$ stand for the optimized parameters after maximization.
In our experiments we simulated QAOA for MaxCut problems with depths $p\in\{1,2,3\}$. For depth $p=1$ the QAOA correlations $b_{ij}^{\mathrm{QAOA}}$ and thus also the expectation value $F(\boldsymbol{\beta},\boldsymbol{\gamma})$ can be expressed analytically as shown in Ref. [30]. In the present work, at $p=1$ we maximize the QAOA expectation value $F(\boldsymbol{\beta},\boldsymbol{\gamma})$ by first applying a rough grid search over the possible parameter space of $\boldsymbol{\beta},\boldsymbol{\gamma}$ before using the gradient-based BFGS algorithm [31] with the best parameters from the grid-search as initial parameters.
For depths $p>1$ we use the Qtensor tensor network library [32, 33, 34] to compute the expectation values and correlations. The basis of the simulation is mapping the quantum circuits that need to be evaluated to tensor networks. This can be done by interpreting a quantum state of $n$ qubits as tensor from $(\mathbb{C}^{2})^{\otimes n}$ and a quantum gate as a tensor with input and output indices for each qubit it acts on [35]. Here, an input index corresponds to the output index of the previous gate. A contraction of the resulting tensor network leads to an exact simulation of the quantum circuit. Because the required computational effort for calculating the result of such a network depends strongly on the order in which the indices are contracted over, it is crucial to optimize the order of contraction [36, 37]. In this work, we use a method of finding a good contraction order based on a line graph representation of the tensor network as introduced in Refs. [33, 32].
Then, we use gradient ascent to find the optimal variational parameters $\boldsymbol{\beta_{\text{opt}}},\boldsymbol{\gamma_{\text{opt}}}$ that maximize $F(\boldsymbol{\beta},\boldsymbol{\gamma})$. To ensure the reliability of the optimization, it is essential to start gradient ascent from suitable initial parameters $\boldsymbol{\beta}_{\text{i}},\boldsymbol{\gamma}_{\text{i}}$ [38]. In this work, we used the so-called fixed-angle conjectures as initial parameters for the tensor network simulations guaranteeing a good performance [39]. Although the conjectures are derived for $k$-regular graphs, we also use them as initial parameters for non-regular graphs with geometries similar to $k$-regular problems. These graphs arise from applying the shrinking steps (introduced in the next subsection) to $k$-regular graphs.

### v2

# Quantum and Classical Correlations in Shrinking Algorithms for Optimization

@@ Abstract
Understanding the benefits of quantum computing for solving combinatorial optimization problems (COPs) remains an open research question. In this work, we extend and analyze algorithms that solve COPs by recursively shrinking them. The algorithms leverage correlations between variables extracted from quantum or classical subroutines to recursively simplify the problem. We compare the performance of the algorithms equipped with correlations from the quantum approximate optimization algorithm (QAOA) as well as the classical linear programming (LP) and semi-definite programming (SDP) relaxations. This allows us to benchmark the utility of QAOA correlations against established classical relaxation algorithms. We apply the recursive algorithm to MaxCut  problem instances with up to a hundred vertices at different graph densities. Our results indicate that LP outperforms all other approaches for low-density instances, while SDP excels for high-density problems. Moreover, the shrinking algorithm proves to be a viable alternative to established methods of rounding LP and SDP relaxations. In addition, the recursive shrinking algorithm outperforms its bare counterparts for all three types of correlations, i.e., LP with spanning tree rounding, the Goemans-Williamson algorithm, and conventional QAOA. While the lowest depth QAOA consistently yields worse results than the SDP, our tensor network experiments show that the performance increases significantly for deeper QAOA circuits.
@@ I Introduction
Combinatorial optimization problems (COPs) are highly relevant in both science and industry. However, they are generally challenging to solve efficiently because of the exponential growth of the solution space with increasing problem size. Traditionally, a variety of classical strategies, including linear programming (LP) [1] and semi-definite programming (SDP) relaxations [2], are applied to find approximate solutions to COPs. Recently, approaches utilizing quantum resources have emerged as promising alternative solution strategies [3]. Among these, the quantum approximate optimization algorithm (QAOA) [4] is particularly interesting because of its universality and adaptive complexity. QAOA is a hybrid quantum-classical variational algorithm designed to find approximate solutions to unconstrained binary optimization problems using noisy-intermediate scale quantum (NISQ) devices and beyond.
@@
It is unclear whether NISQ algorithms can provide improvements over classical methods for solving relevant COPs in industry. In fact, several issues of the near-term variational approaches have been identified, such as barren plateaus [5] and noise of quantum devices [6]. Another restriction of QAOA is due to its locality, i.e., the effect that only qubits separated by less than a certain distance in the graph representation of a problem can interact with each other at a given circuit depth. This property implies limitations on the performance of QAOA [7, 8, 9, 10].
@@
The recursive QAOA (RQAOA) was introduced by Bravyi et al. [7, 11] to overcome the locality-induced limitations of QAOA. The algorithm operates by recursively simplifying the problem based on correlations between variables obtained from QAOA. The two steps of computing the correlations and fixing the variables are executed iteratively until the problem is fully solved. By introducing new connections between previously unlinked variables, RQAOA overcomes the locality of QAOA. Several works have generalized and extended RQAOA. These proposals include using problem-specific update rules [12, 13], analog quantum devices [12] or a shrinking procedure based on the classical calculation of correlations [14]. It has been shown that those recursive algorithms outperform the original QAOA for many problem instances [15, 11, 16]. However, it is not clear whether the improved results are due to the recursive shrinking procedure or the quantum correlations. Here, we aim to give insights into this question.
Contributions
@@
Our study evaluates how different classical and quantum methods for calculating correlations affect the performance of a shrinking algorithm similar to RQAOA. To this end, we introduce new approaches of calculating correlations, beyond the known QAOA, through classical means. The investigated routines for calculating correlations are the classical LP and SDP relaxations, and QAOA. Our numerical experiments on instances of MaxCut allow us to benchmark the utility of quantum correlations against correlations obtained from established classical approximation algorithms. Moreover, the SDP shrinking procedure introduced in this work is a novel recursive approximation algorithm inspired by the Goemans-Williamson approach, specifically designed for MaxCut problems. It outperforms the original Goemans-Williamson algorithm for the randomized problem instances considered in this study. Furthermore, we show that the shrinking algorithm employed here is a viable alternative to traditional rounding routines for LP and SDP relaxations.
@@
The remainder of the paper is organized as follows. In Section II we introduce the MaxCut problem and the different routines (LP, SDP and QAOA) for calculating correlations. In addition, we present the shrinking algorithm that utilizes the calculated correlations. In Section III we analyze the performance of the shrinking algorithm equipped with different correlation sources before discussing the implications of our findings in Section IV. In Section V we conclude by suggesting potential future research directions.
@@ II Methods
In this section, we first introduce the MaxCut problem. Next, we explain how to compute the correlations using the LP, SDP, and QAOA. Finally, we present the shrinking algorithm that uses the correlations to recursively simplify the optimization problem.
@@ II-A MaxCut problem and its encoding
A MaxCut problem instance is defined by a weighted, undirected graph $G=(V,E)$ with vertices $V=\{i\}$, edges $E=\{e\}$ and edge weights $w_{e}$. We denote the number of vertices by $n=|V|$. In MaxCut, we are tasked with finding a subset of nodes that maximizes the weight of edges connecting the chosen node subset and its complement. Formally, we want to find a node partition $W\subseteq V$ such that the edge set $\delta(W)\coloneqq\{ij\in E\mid i\in W,j\in V\setminus W\}$ maximizes its weight defined as $\sum_{e\in\delta(W)}w_{e}$.
Despite its simplicity, MaxCut is a popular example of a COP since any quadratic unconstrained binary optimization (QUBO) problem can be transformed into a MaxCut problem [17, 18]. Furthermore, solving it is NP-hard [19], and finding solutions for dense instances with hundreds of variables can already overstrain state-of-the-art algorithms [20].
The MaxCut problem can be formulated as maximizing an integer quadratic unconstrained cost function $C(\mathbf{x})$ in the form of
@@ II-B Means of computing correlations
The core of our algorithm are correlations between decision variables of a MaxCut problem. Each edge $ij\in E$ in the problem graph is assigned a correlation $b_{ij}\in[-1,1]$. Ideally, we want $b_{ij}$ to represent the correlation between variables in high-quality solutions of the MaxCut problem. In this case, a large negative (positive) correlation between two variables indicates that in high-quality solutions, these two variables mostly take opposite (equal) values. In the language of MaxCut, this translates to an edge predominantly being cut for negative correlations and not cut for positive correlations.
The shrinking algorithm works identically irrespective of the correlations that are used. However, different means of computing correlations might be better suited for different instances. Thus, in this paper, we compare three means of computing correlations: LP and SDP relaxations, and QAOA. All three run in polynomial time using classical or quantum resources.
@@ II-B1 Linear programming (LP)
For a weighted, undirected graph $G=(V,E)$, we define an edge-incidence vector $\mathbf{y}\in\{0,1\}^{|E|}$. Here, a value of $y_{e}=1$ implies that edge $e$ is cut, while $y_{e}=0$ implies that edge $e$ is not cut.
With this notion, we model MaxCut as the integer linear program (cf. [21, 22, 23, 24])
@@
$\begin{split}&y_{e}\in\{0,1\},\ \forall e\in E.\end{split}$ (2d)
Here, the odd cycle inequalites (2b) enforce the number of cut edges in any cycle to be even, which is sufficient for $y_{e}$ to define a cut. While the MaxCut  problem defined by Eqs. (2a), (2b), (LABEL:eq:_odd-cycle_d) cannot be solved efficiently in general, the linear relaxation defined by Eqs. (2a)–(LABEL:eq:_odd-cycle_c) can be solved in polynomial time and yields an upper bound on the maximum cut [20]. This also means that, if the relaxation returns an integer solution, the problem is solved to optimality.
Analogously to Ref. [14], we compute correlations via the affine function
$b_{e}^{\mathrm{LP}}\coloneqq 1-2\tilde{y}_{e}\in[-1,1]$ (3)
where $\tilde{\mathbf{y}}\in[0,1]^{|E|}$ is an optimal solution to the linear program Eqs. (2a)–(LABEL:eq:_odd-cycle_c). Thus, we assign strong correlations to edges whose LP value is close to either 00 or $1$.
@@ II-B2 Semi-definite programming (SDP)
The calculation of semi-definite programming correlations is inspired by the seminal approximation algorithm introduced by Goemans and Williamson [2]. The main idea underlying the Goemans-Williamson (GW) algorithm is that we replace the NP-hard integer quadratic problem formulation of MaxCut as introduced in Eq. (1) by a relaxed version of the problem. The relaxation admits a larger solution space that contains all possible solutions to the integer problem. Thus, the optimal solution of the relaxed problem upper bounds the best solution of the integer model by design. Concretely, we replace integer values with multi-dimensional vectors $\mathbf{v}_{i}\in\mathbb{R}^{n}$ with $\norm{\mathbf{v}_{i}}=1$, i.e., we expand the search space to the $(n-1)$-dimensional unit sphere $S_{n-1}$. Using this, we can relax the maximization of Eq. (1) to
@@
where $L\coloneqq\text{diag}(A\mathbf{{e}})-A$ is the Laplacian of the problem graph, $A$ denotes the adjacency matrix, and $\mathbf{{e}}$ stands for the vector of all ones. For details see Refs. [2, 25]. Here, $\langle a,b\rangle=\text{tr}(ab^{\text{T}})$ represents the Frobenius inner product of two matrices $a$, $b\in\mathbb{R}^{n\times m}$. This generic semi-definite program can be solved in polynomial time by various solvers such as cvxopt, see, e.g., [26, 27, 25].
@@
From the vectors $\{\mathbf{v}_{i}\}$, we obtain the correlation of edge $ij\in E$ by evaluating the dot product
@@
which measures the degree of (anti-)alignment of the vectors. Thus, we assign a large absolute correlation if the vectors are approximately parallel or anti-parallel aligned. The SDP correlations (6) are motivated by the integer quadratic model (1), where an edge is cut if $x_{i}x_{j}=-1$, and not cut if $x_{i}x_{j}=+1$. However, in contrast to the integer model (1), the SDP correlations can take continuous values between $-1$ and $+1$. Furthermore, we can compute these correlations efficiently.
@@
Alternatively, we can obtain correlations by leveraging the hyperplane rounding procedure from the GW algorithm [2]. The rounding procedure maps an optimal solution of (4) to a MaxCut solution. To this end, we first choose a random hyperplane, defined by a normal vector $\mathbf{r}\in S_{n-1}$. Next, we divide the vertices $V$ into two partitions $S\subseteq V$ and $\overline{S}=V\,\backslash\,S$ using the criterion
$S=\{i\in V|\ \mathbf{v}_{i}\cdot\mathbf{r}\geq 0\}.$ (7)
Since we can efficiently evaluate the objective value of a given candidate cut, we may try several (polynomially many) random hyperplanes and return the best solution found. Ref. [2] shows that the cuts obtained this way are guaranteed to achieve an approximation ratio of at least $0.878$ on average.
@@
The GW correlations are designed to combine both the global information about the partition of the vertices and the local information about the (anti-)parallel alignment of the vectors.
@@ II-B3 QAOA
Finally, we compute correlations using QAOA. Here, we choose QAOA is a representative for various quantum routines. Therefore, we remark that QAOA can readily be replaced by other quantum approaches, such as warm-start QAOA [28] or Quantum Annealing [29]. We choose it because of its wide spread within research to achieve a proof of concept. The goal of QAOA is to find an $\mathbf{x}\in\{-1,1\}^{n}$ that maximizes the cost function $C:\{-1,1\}^{n}\rightarrow\mathbb{R}$ of an arbitrary integer quadratic unconstrained optimization problem. Starting from the uniform superposition state $\ket{+}$ defined by
@@
is maximized. Thus, we aim at preparing a superposition of high quality candidate solutions.
It follows from Eq. (1) that, in the case of MaxCut , the cost Hamiltonian $H_{C}$ is given by
@@
where $I$ is the identity operator and $Z_{i}$ is the Pauli-Z operator acting on qubit $i$. After maximizing the expectation value $F(\boldsymbol{\beta},\boldsymbol{\gamma})$, we compute correlations $b_{ij}$ between nodes $i$ and $j$ via
@@
where $\boldsymbol{\beta_{\text{opt}}}$ and $\boldsymbol{\gamma_{\text{opt}}}$ are the optimized parameters.
In our experiments, we simulate QAOA with depths $p\in\{1,2,3\}$. Ref. [30] derives analytcial expressions for both the QAOA correlations $b_{ij}^{\mathrm{QAOA}}$ and the expectation value $F(\boldsymbol{\beta},\boldsymbol{\gamma})$ in the case $p=1$. In the present work, we maximize the QAOA expectation value $F(\boldsymbol{\beta},\boldsymbol{\gamma})$ for $p=1$ by first applying a rough grid search over the parameter space of $\boldsymbol{\beta},\boldsymbol{\gamma}$. Then, we apply the gradient-based BFGS algorithm [31] with the best parameters from the grid-search as initial parameters.
For depths $p>1$, we use the Qtensor tensor network library [32, 33, 34] to compute the expectation values and correlations. The library maps the quantum circuits we want to evaluate to tensor networks. This is done by interpreting a quantum state of $n$ qubits as a tensor from $(\mathbb{C}^{2})^{\otimes n}$ and a quantum gate as a tensor with input and output indices for each qubit it acts on [35]. Here, an input index corresponds to the output index of the previous gate. A contraction of the resulting tensor network leads to an exact simulation of the quantum circuit. Because the required computational effort depends strongly on the order in which the indices are contracted, it is crucial to optimize the order of contraction [36, 37]. In this work, we use a method of finding a good contraction order based on a line graph representation of the tensor network as introduced in Refs. [33, 32].
We use gradient ascent to find the optimal variational parameters $\boldsymbol{\beta_{\text{opt}}},\boldsymbol{\gamma_{\text{opt}}}$ that maximize $F(\boldsymbol{\beta},\boldsymbol{\gamma})$. To ensure the reliability of the optimization, it is essential to start gradient ascent from suitable initial parameters $\boldsymbol{\beta}_{\text{i}},\boldsymbol{\gamma}_{\text{i}}$ [38]. In this work, we use the so-called fixed-angle conjectures as initial parameters for the tensor network simulations guaranteeing a good performance [39]. Although the conjectures are derived for $k$-regular graphs, we also use them as initial parameters for non-regular graphs with geometries similar to $k$-regular problems. These graphs arise from applying the shrinking steps (introduced in the next subsection) to $k$-regular graphs. Our experiments show that QAOA, using parameters from the fixed-angle conjecture as initialization, produces output states with energies that are at least as good as those achieved with established parameter initializations such as using transition states [40] or interpolation methods [41].
We would like to stress that even though QAOA can be run in polynomial time on a quantum computer (for fixed $p$), it requires significant resources to simulate the algorithm using tensor networks. Moreover, the time required to simulate QAOA classically depends strongly on the MaxCut problem at hand. To keep computations tractable, we restrict the simulation to 3-regular graphs.